# A Vision-Language Pre-training Model-Guided Approach for Mitigating Backdoor Attacks in Federated Learning

### v1

@@ Abstract
Existing backdoor defense methods in Federated Learning (FL) rely on the assumption of homogeneous client data distributions or the availability of a clean serve dataset, which limits the practicality and effectiveness. Defending against backdoor attacks under heterogeneous client data distributions while preserving model performance remains a significant challenge. In this paper, we propose a FL backdoor defense framework named CLIP-Fed, which leverages the zero-shot learning capabilities of vision-language pre-training models. By integrating both pre-aggregation and post-aggregation defense strategies, CLIP-Fed overcomes the limitations of Non-IID imposed on defense effectiveness. To address privacy concerns and enhance the coverage of the dataset against diverse triggers, we construct and augment the server dataset using the multimodal large language model and frequency analysis without any client samples. To address class prototype deviations caused by backdoor samples and eliminate the correlation between trigger patterns and target labels, CLIP-Fed aligns the knowledge of the global model and CLIP on the augmented dataset using prototype contrastive loss and Kullback-Leibler divergence. Extensive experiments on representative datasets validate the effectiveness of CLIP-Fed. Compared to state-of-the-art methods, CLIP-Fed achieves an average reduction in ASR, i.e., 2.03% on CIFAR-10 and 1.35% on CIFAR-10-LT, while improving average MA by 7.92% and 0.48%, respectively.
Codehttps://anonymous.4open.science/r/CLIP-Fed
Introduction
Figure 1: An overall framework of CLIP-Fed. We utilize the prior knowledge and cross-modal visual-language representation capabilities of CLIP to purify backdoor.
@@ Introduction
The robustness of Federated Learning (FL) (Han et al. 2023; Fang, Ye, and Du 2025) has become a growing concern, particularly under adversarial settings. The decentralized training process makes FL vulnerable to backdoor attacks (Zhuang et al. 2024), as malicious clients can poison local training data to induce the global model and associate specific trigger patterns (e.g., pixel patches or watermarks) with target labels. Due to FL’s distributed nature, the backdoor attacks become more stealthy and cause significantly more severe damage than in centralized settings (Liao et al. 2025; Lyu et al. 2025).
@@
Existing defense strategies can be classified into three main categories. Specifically, the first type distinguishes benign and malicious clients by clustering model parameters (Nguyen et al. 2022; Mu et al. 2024; Fereidooni et al. 2024), while highly relies on the homogeneous data distribution of FL. The second depends on modified aggregation strategies such as weighted aggregation, median-based aggregation, and rank-based aggregation to achieve robust FL (Han et al. 2023; Cao et al. 2021; Yin et al. 2018). However, malicious clients can adjust attack strategies to circumvent these rules, such as the screening of poisoning parameters and hierarchical poisoning mechanisms (Zhuang et al. 2024). The third kind of methods use differential privacy to reduce the impact of backdoor attacks, but task performance is generally negatively affected (Nguyen et al. 2022; Bagdasaryan et al. 2020). Thus, most existing methods are built on overly rigid assumptions of accessing to a clean labeled validation set or homogeneous data distribution across clients, which limit their real-world utility. Moreover, existing defense methods overlook the issue of class prototype deviations caused by the presence of backdoor samples, which leads to a degradation in model performance during the backdoor defense process.
@@
To address above challenges, we propose to leverage the rich prior knowledge and cross-modal visual-language representation capabilities of Vision-Language Pretraining models (VLPs) to defend the global model against backdoor attacks, as shown in Figure 1. CLIP (Radford et al. 2021) is a representative VLP that uses contrastive learning, and its powerful zero-shot learning capability makes it a potential tool for purifying backdoor patterns in the global model. First, we integrate both pre- and post-aggregation defense strategies to eliminate reliance on homogeneous client data distributions. Specifically, pre-aggregation model clustering filters out malicious updates, while post-aggregation purification uses a global model adjustment to enhance robustness under Non-IID conditions. To further address privacy concerns, we construct a server dataset by combining Multimodal Large Language Models (MLLMs) and frequency analysis without any client samples, which balances privacy protection, coverage of diverse backdoor triggers, and the impact of image augmentation on model utility. Finally, by using the enhanced server dataset under CLIP’s semantic guidance via prototype contrastive learning, we rectify the global model’s feature extractor to correct class prototype deviations caused by backdoor samples, while decoupling the strong correlations between trigger patterns and target labels. To this end, we align the logits distributions of CLIP and the global model through knowledge transfer to further remove residual backdoor associations.
@@
The main contributions are summarized as follows. (1) We propose ̵CLIP-Fed, the first ̵VLP-guided FL backdoor defense framework, which uses the cross-modal representation capabilities of VLPs to mitigate backdoor attacks. CLIP-Fed combines pre-aggregation model clustering and post-aggregation backdoor purification to jointly enhance defense robustness, reducing dependence on homogeneous client data distributions. (2) To correct ̵class prototype deviations induced by backdoor samples, we propose a ̵semantically supervised global model rectification method using CLIP, which employs a ̵prototype contrastive loss and ̵Kullback-Leibler (KL) divergence to align the global model features with CLIP’s embeddings in an enhanced dataset, thereby ̵eliminating correlations between backdoor triggers and target labels. (3) Extensive evaluations on ̵three datasets and ̵four backdoor attack types demonstrate CLIP-Fed’s superiority. Compared to state-of-the-art methods, CLIP-Fed achieves an average reduction in ASR, i.e., 2.03% on CIFAR-10 and 1.35% on CIFAR-10-LT, while improving average MA by 7.92% and 0.48%, respectively. Additionally, CLIP-Fed maintains ̵client-side efficiency without extra computational overhead beyond local training.
Related Work
@@
Backdoor Defense in FL. To mitigate the backdoor attacks, for example, Auror (Shen, Tople, and Saxena 2016), Flame (Nguyen et al. 2022), FedDMC (Mu et al. 2024), and FreqFed (Fereidooni et al. 2024) tried model clustering to eliminate malicious backdoors while ensuring the performance of the aggregated model. The effectiveness of such methods depended on an assumption of homogeneous client data distribution, such that the performance under Non-IID conditions is limited. Other methods, e.g., BRCA (Zhai et al. 2021) and FLTrust (Cao et al. 2021), assign trust scores to clients using a small clean validation set hosted on the server to detect and filter malicious contributions. While effective in some cases, this paradigm contradicts the decentralized and privacy-preserving nature of FL by requiring centralized access to client data. Recent work has explored contrastive learning as a defense mechanism. Through contrastive learning (Radford et al. 2021), FLPurifier (Zhang et al. 2024) severs latent correlations between target labels and trigger patterns, thereby neutralizing potential backdoor artifacts during model training phases. Although this strategy improves model robustness, it increases the computational burden on the client and makes it less suitable for resource-constrained environments. Overall, achieving effective backdoor defense under Non-IID conditions while ensuring privacy and avoiding increasing client-side overhead remains a critical challenge. Our scheme proposes a server-side dataset augmentation strategy based on MLLMs (Lee et al. 2024) and frequency analysis, which enhances the dataset’s coverage against diverse backdoor triggers without accessing client data. Additionally, our scheme integrates a global purification mechanism combining prototype contrastive learning and knowledge distillation, such that effective suppression of backdoor patterns in Non-IID FL can be achieved.
@@
Figure 2: The framework of the proposed CLIP-Fed. CLIP-Fed contains four modules: Data Augmentation for Server Dataset aims to build the server dataset, Dynamic Model Filtering by Clustering aims to filter out malicious models before aggregation, Feature Rectification via Prototype Alignment and Global Model Knowledge Transfer aim to purify the backdoor.
@@
Vision-Language Pre-train Model. VLPs (Radford et al. 2021; Li et al. 2022; Liu et al. 2023) have made significant strides across various downstream tasks such as image classification, visual question answering, and cross-modal retrieval. CLIP (Radford et al. 2021) is a representative VLP that uses contrastive learning to capture semantic relationships between images and texts, enabling both to be mapped into a shared feature space. Recent studies have also explored integrating contrastive learning into FL to address the challenges of Non-IID client data. MOON (Li, He, and Song 2021) proposes a model-contrastive FL framework that utilizes inter-client representation similarity to improve model generalization under heterogeneous data distributions. In addition, the strong semantic understanding of VLPs has shown promise in handling long-tailed distributions in FL (Shi et al. 2024), as VLPs can capture rich visual concepts even from underrepresented classes. However, while such capabilities indicate that VLPs are well suited for learning robust and semantically meaningful representations, their potential for enhancing security in FL, particularly in defending against backdoor attacks, remains underexplored. To the best of our knowledge, our work is the first to use the VLPs for backdoor defense in FL. Leveraging the cross-modal representation capacity of VLPs, we align the global model’s visual features with those of a VLP to detect and eliminate backdoor patterns, establishing a novel and practical defense paradigm.
Methodology
Problem Formulation
@@
In the FL setup with one server and $m$ clients, each client $i$ holds a local dataset $D_{i}$ and a model $w_{i}=(f_{i},c_{i})$, where $f_{i}$ is the feature extractor and $c_{i}$ is the classifier. The server maintains a global model $G_{s}=(f_{s},c_{s})$. For an input $x$, the prediction is obtained via $p=c(f(x))$. Malicious clients conduct backdoor attacks by embedding triggers in training samples and uploading poisoned models. The manipulated models aim to cause the global model to misclassify triggered inputs as the target label $y^{\prime}$. The Attack Success Rate (ASR) is defined as:
@@
$\mathrm{ASR}=\underset{\begin{subarray}{c}(x,y)\sim D\\ y\neq y^{\prime}\end{subarray}}{\mathbb{E}}\left[\operatorname{Pr}\left(\Phi(t(x))=y^{\prime}\right)\right],$ (1)
where $D$ is the dataset for the FL task, $y$ is the label for input $x$, $y^{\prime}$ is the target label, and $t(\cdot)$ is the backdoor trigger function. The framework of CLIP-Fed is shown in Figure 2. CLIP-Fed aims to suppress backdoor and reduce ASR.
Data Augmentation for Server Dataset
Generation of Image-Text Samples
@@
In CLIP-Fed, to preserve client data privacy, the server cannot directly utilize client datasets for contrastive learning and knowledge distillation. MLLMs can be utilized to perform server-side dataset generation and augmentation without accessing client data (Lee et al. 2024; Pan et al. 2025). Unlike traditional methods that rely on geometric transformations or noise injection, our scheme semantically expands candidate labels to guide the image generation. For instance, combining object descriptors with contextual modifiers such as “a small airplane beneath a cloudy sky” or “a frog partially obscured by leaves” naturally introduces variations in scale, occlusion, and background complexity, which inherently influence visual perception. The synthesized images are then adaptively resized and normalized to match the statistical distribution of the target dataset, such that the alignment with the original feature space is guaranteed.
Frequency Sensitivity-guided Image Augmentation
@@
To enhance the robustness of backdoor purification to triggers, we add perturbations to image samples of the server dataset to simulate images with triggers while preserving label information. Specifically, we first add visible triggers to a clean image $x$ to obtain $\bar{x}$. Then, we apply the Discrete Cosine Transform (DCT) to both $x$ and $\bar{x}$ and divide the spectra into multiple frequency bands. We then identify the frequency intervals most sensitive to triggers by comparing the similarities between images $x$ and $\bar{x}$ in different frequency domains. We assume that frequency intervals with the lowest similarity have the highest sensitivity to triggers. The DCT spectrogram shows low-frequency coefficients clustered in the upper-left and high-frequency components in the lower-right, symmetric about the main diagonal. We segment the spectrogram into discrete frequency bands ordered from low to high, aligned orthogonally to the diagonal. Specifically, the matrix is unevenly divided into $K$ segments along the diagonal, with the distance from the $k$-th division point to the top-left corner set to $1/2^{(K-k)}$ of the diagonal length. We then extend these $K$ frequency bands into $K$ independent spectrograms, represented as $\left(F_{1}(u,v),\ldots,F_{K}(u,v)\right)$. We assess the distance between the spectrogram coefficients in each frequency segment using Mean Squared Error (MSE). Frequency regions with higher MSE values exhibit lower similarity, indicating higher sensitivity to pixel-block triggers. The frequency band with the highest MSE is denoted as $F_{high}(u,v)$, where we inject Gaussian noise perturbations into the original image. The image is then reconstructed into the spatial domain via IDCT, yielding an enhanced server dataset image with noise added only in the sensitive $F_{high}$ region. Perturbing only the most sensitive frequency bands better simulates backdoor triggers while minimizing impact on model utility.
Dynamic Model Filtering by Clustering

### v2

@@ Abstract.
Defending backdoor attacks in Federated Learning (FL) under heterogeneous client data distributions encounters limitations balancing effectiveness and privacy-preserving, while most existing methods highly rely on the assumption of homogeneous client data distributions or the availability of a clean serve dataset. In this paper, we propose an FL backdoor defense framework, named CLIP-Fed, that utilizes the zero-shot learning capabilities of vision-language pre-training models. Our scheme overcomes the limitations of Non-IID imposed on defense effectiveness by integrating pre-aggregation and post-aggregation defense strategies. CLIP-Fed aligns the knowledge of the global model and CLIP on the augmented dataset using prototype contrastive loss and Kullback-Leibler divergence, so that class prototype deviations caused by backdoor samples are ensured and the correlation between trigger patterns and target labels is eliminated. In order to balance privacy-preserving and coverage enhancement of the dataset against diverse triggers, we further construct and augment the server dataset via using the multimodal large language model and frequency analysis without any client samples. Extensive experiments on representative datasets evidence the effectiveness of CLIP-Fed. Comparing to other existing methods, CLIP-Fed achieves an average reduction in Attack Success Rate, i.e., 2.03% on CIFAR-10 and 1.35% on CIFAR-10-LT, while improving average Main Task Accuracy by 7.92% and 0.48%, respectively. Our codes are available at https://anonymous.4open.science/r/CLIP-Fed.
Federated Learning, Backdoor Defense, Vision-Language Pre-training Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Distributed systems security
1. Introduction
@@
The robustness of Federated Learning (FL) (Han et al., 2023; Wang et al., 2025a; Fang et al., 2025; Gai et al., 2025; Miao et al., 2024; Wang et al., 2025b) is a growing concern since the distributed training process makes FL vulnerable to backdoor attacks (Zhuang et al., 2024; Shen et al., 2025; Zhang et al., 2024a; Li et al., 2025), i.e., malicious clients can poison local training data to induce the global model and associate specific trigger patterns (e.g., pixel patches or watermarks) with target labels. Backdoor attacks are more difficult to detect in FL’s distributed computing paradigm than in a centralized computing environment (Liao et al., 2025; Lyu et al., 2025).
Figure 1. An overall framework of CLIP-Fed. We utilize the prior knowledge and cross-modal visual-language representation capabilities of CLIP to purify backdoor.
@@
However, existing defense strategies can hardly balance the contradiction between privacy protection and effectiveness, due to reliance of a common assumption on the clean labeled validation set or homogeneous data distribution across clients. Existing methods also overlook the issue of class prototype deviations caused by the presence of backdoor samples, which results in a degradation in model performance during the backdoor defense process. To be specific, the first type of existing methods distinguishes benign and malicious clients by clustering model parameters (Nguyen et al., 2022; Mu et al., 2024; Fereidooni et al., 2024), while it is highly dependent on the homogeneous data distribution of FL. The second depends on modified aggregation strategies to achieve robust FL, e.g., weighted, median-based, and rank-based aggregations, but malicious clients can adjust attack strategies to circumvent these rules, such as the screening of poisoning parameters and hierarchical poisoning mechanisms (Han et al., 2023; Cao et al., 2021; Yin et al., 2018; Zhuang et al., 2024). Finally, another strategy applies differential privacy to reduce the impact of backdoor attacks, causing negative impact on performance (Nguyen et al., 2022; Bagdasaryan et al., 2020).
@@
To address challenges above, we propose an FL backdoor defense framework, called CLIP-Fed, that uses rich prior knowledge and cross-modal visual-language representation capabilities of Vision-Language Pretraining models (VLPs) to defend the global model against backdoor attacks, as shown in Figure 1. CLIP (Radford et al., 2021) is a representative VLP possessing zero-shot learning capability such that it can be used for purifying backdoor patterns in the global model. First, we integrate pre- and post-aggregation defense strategies to eliminate reliance on homogeneous client data distributions. Specifically, pre-aggregation model clustering filters out malicious updates, and post-aggregation purification uses a global model adjustment to enhance robustness under Non-IID conditions. To further address privacy concerns, we construct a server dataset by combining Multimodal Large Language Models (MLLMs) and frequency analysis without any client samples. Finally, by using the enhanced server dataset under CLIP’s semantic guidance via prototype contrastive learning, we rectify the global model’s feature extractor to correct class prototype deviations that are caused by backdoor samples, while decoupling strong correlations between trigger patterns and target labels. To this end, we align the logits distributions of CLIP and the global model through knowledge transfer to further remove residual backdoor associations.
@@
Main contributions of our work are summarized as follows. (1) We propose ̵CLIP-Fed, the first ̵VLP-guided FL backdoor defense framework, which uses the cross-modal representation capabilities of VLPs to mitigate backdoor attacks. CLIP-Fed combines pre-aggregation model clustering and post-aggregation backdoor purification to jointly enhance defense robustness, reducing dependence on homogeneous client data distributions. (2) To correct ̵class prototype deviations induced by backdoor samples, we propose a ̵semantically supervised global model rectification method via CLIP, which employs a ̵prototype contrastive loss and ̵Kullback-Leibler (KL) divergence to align global model features with CLIP’s embeddings in an enhanced dataset, thereby ̵eliminating correlations between backdoor triggers and target labels. (3) Extensive evaluations on ̵three datasets and ̵four backdoor attack types demonstrate CLIP-Fed’s superiority. Compared to existing methods, CLIP-Fed achieves an average reduction in ASR, i.e., 2.03% on CIFAR-10 and 1.35% on CIFAR-10-LT, while improving the average MA by 7.92% and 0.48%, respectively. CLIP-Fed maintains ̵client-side efficiency without extra computational overhead beyond local training.
2. Related Work
@@
Backdoor Defense in FL. To mitigate the backdoor attacks, for example, Auror (Shen et al., 2016), Flame (Nguyen et al., 2022), FedDMC (Mu et al., 2024), and FreqFed (Fereidooni et al., 2024) tried model clustering to eliminate malicious backdoors while ensuring the performance of the aggregated model. The effectiveness of such methods depended on an assumption of homogeneous client data distribution, such that practicability was limited under Non-IID conditions. Other methods, e.g., BRCA (Zhai et al., 2021) and FLTrust (Cao et al., 2021), assigned trust scores to clients using a small clean validation set hosted on the server to detect and filter malicious contributions. While effective in some cases, this paradigm contradicted the decentralized and privacy-preserving nature of FL by requiring centralized access to client data. Recent work has explored contrastive learning as a defense mechanism. Through contrastive learning (Radford et al., 2021), FLPurifier (Zhang et al., 2024b) eliminated latent correlations between target labels and trigger patterns, thus neutralizing potential backdoor artifacts during model training phases. Although this strategy improved model robustness, it increased the computational burden on the client and made it less suitable for resource-constrained environments. Overall, achieving effective backdoor defense under Non-IID conditions while ensuring privacy and avoiding increasing client-side overhead remained a critical challenge. Our scheme is a server-side dataset augmentation strategy based on MLLMs (Lee et al., 2024) and frequency analysis, which enhances the dataset’s coverage against diverse backdoor triggers without accessing client data. Additionally, our scheme achieves effective suppression of backdoor patterns in Non-IID FL via integrating a global purification mechanism combining prototype contrastive learning and knowledge distillation.
@@
Figure 2. The framework of the proposed CLIP-Fed. CLIP-Fed contains four modules: Data Augmentation for Server Dataset aims to build the server dataset, Dynamic Model Filtering by Clustering aims to filter out malicious models before aggregation, Feature Rectification via Prototype Alignment and Global Model Knowledge Transfer aim to purify the backdoor.
@@
Vision-Language Pre-train Model. VLPs (Radford et al., 2021; Li et al., 2022; Liu et al., 2023) had made significant strides across various downstream tasks, e.g., image classification, visual question answering, and cross-modal retrieval. CLIP (Radford et al., 2021) was a typical VLP that used contrastive learning to capture semantic relationships between images and texts, enabling both to be mapped into a shared feature space. Recent studies tried to integrate contrastive learning into FL to address the challenges of Non-IID client data. MOON (Li et al., 2021) was a model-contrastive FL framework that utilized inter-client representation similarity to improve model generalization under heterogeneous data distributions. In addition, the strong semantic understanding of VLPs had shown promise in handling long-tailed distributions in FL (Shi et al., 2024), as VLPs could capture rich visual concepts even from underrepresented classes. However, while such capabilities indicated that VLPs were well suited for learning robust and semantically meaningful representations, the potential to improve security in FL remained underexplored, particularly in defending against backdoor attacks. To the best of our knowledge, our work is the first to use VLPs for backdoor defense in FL. Using the cross-modal representation capacity of VLPs, we align the global model’s visual features with those of a VLP to detect and eliminate backdoor patterns, establishing a novel and practical defense paradigm.
3. Methodology
3.1. Problem Formulation
@@
In the FL setup with one server and $m$ clients, each client $i$ holds a local dataset $D_{i}$ and a model $w_{i}=(f_{i},c_{i})$, where $f_{i}$ is the feature extractor and $c_{i}$ is the classifier. The server maintains a global model $G_{s}=(f_{s},c_{s})$. For an input $x$, the prediction is obtained via $p=c(f(x))$. Malicious clients conduct backdoor attacks by embedding triggers in training samples and uploading poisoned models. The manipulated models aim to cause the global model to misclassify triggered inputs as the target label $y^{\prime}$. Attack Success Rate (ASR) is defined by Eq. (1), where $D$ is the dataset for the FL task, $y$ is the label for input $x$, $y^{\prime}$ is the target label, and $t(\cdot)$ is the backdoor trigger function. The framework of CLIP-Fed is shown in Figure 2. CLIP-Fed aims to suppress backdoor and reduce ASR. The algorithm flow of CLIP-Fed is shown in Alg. 1.
@@
(1) $\mathrm{ASR}=\underset{\begin{subarray}{c}(x,y)\sim D\\ y\neq y^{\prime}\end{subarray}}{\mathbb{E}}\left[\operatorname{Pr}\left(\Phi(t(x))=y^{\prime}\right)\right].$
Algorithm 1 Training Process of CLIP-Fed
0: Initial global model $G_{1}(f_{s}^{1},c_{s}^{1})$, number of rounds $T$, NN clients
0: Global model $G_{T+1}(f_{s}^{T+1},c_{s}^{T+1})$
[…]
3:for each training iteration $t$ $\in$ $[1,T]$ do
4:  for each client $i\in C_{t}$ do
5:   $f_{i}^{t},c_{i}^{t}\leftarrow{\text{Local Training}}(G_{t},i)$
6:  end for
7:  $S_{t}\leftarrow{\text{HDBSCAN Clustering}(\{f_{1}^{t},c_{1}^{t}\},\dots,\{f_{i}^{t},c_{i}^{t}\})}$
8:  $f_{s}^{t+1}\leftarrow\sum_{i\in S_{t}}\frac{D_{i}}{D_{S_{t}}}f_{i}^{t},\ c_{s}^{t+1}\leftarrow\sum_{i\in S_{t}}\frac{D_{i}}{D_{S_{t}}}c_{i}^{t}$
9:  Feature extractor $f_{s}^{t+1}$ rectification by prototype contrastive loss
10:  Global model $(f_{s}^{t+1},c_{s}^{t+1})$ knowledge transfer by KL divergence
11:  $G_{t+1}\leftarrow\{f_{s}^{t+1},c_{s}^{t+1}\}$
12:end for
13:return $G_{T+1}$
3.2. Data Augmentation for Server Dataset
3.2.1. Generation of Image-Text Samples
@@
In CLIP-Fed, the server cannot directly utilize client datasets for contrastive learning and knowledge distillation for the purpose of privacy-preserving. MLLMs are used to generate and augment server-side datasets without accessing client data (Lee et al., 2024; Pan et al., 2025). Dissimilar to other existing methods that rely on geometric transformation or noise injection, our scheme semantically expands candidate labels to guide image generation. For instance, combining object descriptors with contextual modifiers, e.g., “a small airplane beneath a cloudy sky” or “a frog partially obscured by leaves”, naturally introduces variations in scale, occlusion, and background complexity, thereby inherently influencing visual perception. The synthesized images are then adaptively resized and normalized to match the statistical distribution of the target dataset, thus the alignment with the original feature space is guaranteed.
3.2.2. Frequency Sensitivity-guided Image Augmentation
@@
To enhance the robustness of backdoor purification against triggers, we add perturbations to image samples of the server dataset to simulate images with triggers while preserving label information. Specifically, we first add visible triggers to a clean image $x$ to obtain $\bar{x}$, then apply the Discrete Cosine Transform (DCT) to both $x$ and $\bar{x}$ and divide their spectra into multiple frequency bands. By comparing the similarities between the images $x$ and $\bar{x}$ in different domains, we identify the frequency intervals that are the most sensitive to triggers. We assume that frequency intervals with the lowest similarity exhibit the highest sensitivity to triggers. The DCT spectrogram displays low-frequency coefficients clustered in the upper-left and high-frequency components in the lower-right, exhibiting symmetric about the main diagonal. We segment the spectrogram into discrete frequency bands ordered from low to high, aligned orthogonally to the diagonal. Specifically, the matrix is unevenly divided into $K$ segments along the diagonal, with the distance from the $k$-th division point to the top-left corner set to $1/2^{(K-k)}$ of the diagonal length. Then, $K$ frequency bands are extended into $K$ independent spectrograms, represented as $\left(F_{1}(u,v),\ldots,F_{K}(u,v)\right)$. We evaluate the distance between the spectrogram coefficients in each frequency segment using Mean Squared Error (MSE). Frequency regions with higher MSE values demonstrate lower similarity, indicating higher sensitivity to pixel-block triggers. The frequency band with the highest MSE is denoted as $F_{high}(u,v)$, where Gaussian noise perturbations are injected into the original image. The image is then reconstructed into the spatial domain via IDCT, resulting in an enhanced server dataset image with noise added only in the sensitive $F_{high}$ region. Perturbing only the most sensitive frequency bands better simulates backdoor triggers while preserving model utility.
3.3. Dynamic Model Filtering by Clustering